•Corum Inc. expects to earn $120 million per year in perpetuity if it does not undertake any new projects. The firm has an opportunity to invest $15 million today, $10 million in one year and $5 million in two years. The cash inflow will grow at a rate of g thereafter. The firm has 5 million shares of common stock outstanding, and the required rate of return on the stock is 15 percent. Ignore taxes and depreciation.   

  • –What is the stock price if the firm does not make the new investment?

  • –What is the impact of new investment on the stock price if g is equal to 2 percent? What is the new stock price?

  • –At what constant growth rate (g) would the new investment just break even?
Answer #1

Without investment

Cash flows per year = $ 120 Million

Required rate of return on stock = 15%

Value of stock = 120/0.15 = $ 800 M

Share price = 800 M/5 M = $ 160 per share

With Investment

Initial Investment, CF0 = -15 M

Cash flow 1 CF1 = 10 M

Cash flow 2, CF2 = 5 M

Terminal Value of cash flows = 5*(1+2%)/(15%-2%) = 39.23 M

NPV of cash flows = -15 + 10/1.15 + 5/(1.15)^2 + 39.23/(1.15)^2 = 27.13 M

Increase in share price = 27.13 M/5 M = $ 5.42 per share

New Stock Price = 160 + 5.42 = $ 165.42